CN2(AS4809) backbone is currently unstable.
The packet loss rate might higher than normal. DMIT NOC has already reached CTGNOC about this situation.
The packet loss rate might higher than normal. DMIT NOC has already reached CTGNOC about this situation.
We put an inappropriate filter policy on the AS4809 interface which lead to unable to trace the routing from DMIT to China.
The inappropriate policy has been removed.
The inappropriate policy has been removed.
LAX management network maintenance.
The panel for LAX will show "node connection failure."
It doesn't affect the actual services.
The panel for LAX will show "node connection failure."
It doesn't affect the actual services.
/var/log/DMIT-NOC.log
LAX management network maintenance. The panel for LAX will show "node connection failure." It doesn't affect the actual services.
Caused by ZenLayer network failure. Resolved.
We'll use ourselves' network in all locations for major management networks. The datacenter's network and LTE will work as redundancy.
We'll use ourselves' network in all locations for major management networks. The datacenter's network and LTE will work as redundancy.
Cloudflare backbone failure leads to high packet loss of the DMIT sPro network.
/var/log/DMIT-NOC.log
Cloudflare backbone failure leads to high packet loss of the DMIT sPro network.
We don't have any response from Cloudflare. Detour it manually to relieve this situation.
This is a long-term promotion code for Lite products for new order only.
Lite-Semi-Annually-Recur-20OFF
Lite-Annually-Recur-30OFF
It’s for annually and semi-annually orders with LITE profile only.
The TINY has been excluded.
Lite is not recommended for China access. Please order in monthly if you think the lite is good for you (at this time.)
Lite-Semi-Annually-Recur-20OFF
Lite-Annually-Recur-30OFF
It’s for annually and semi-annually orders with LITE profile only.
The TINY has been excluded.
Lite is not recommended for China access. Please order in monthly if you think the lite is good for you (at this time.)
CMI:
The cable fault leads to high latency between Shanghai - Tokyo
> Thanks for waiting. As checked, due to submarine cable maintenance, our upstream backbone is affected till 15 Dec. Currently, it is already the best path and no room to optimize.
The cable fault leads to high latency between Shanghai - Tokyo
> Thanks for waiting. As checked, due to submarine cable maintenance, our upstream backbone is affected till 15 Dec. Currently, it is already the best path and no room to optimize.
/var/log/DMIT-NOC.log
Tokyo Pro AS4837 issue notice: CTGnet is still working on us for the optimization. Currently, CU didn't accept our announcement via AS23764. Two timeline: - Next week: AS10099 go live. - Unknow: CTGnet and CU fixed their AS-PATH: 4837 4134 23764 54574.
After 2 months.
Finally, 4837 4134 4809(23764) 54574.
Finally, 4837 4134 4809(23764) 54574.
We are going to out of resources in Tokyo. Due to the global chip shortage, the costs of servers have been doubled.
Our team decided to buy new servers after 2021.
As we estimate the price will be normal between Feb ~ Mar 2022
Before that, once our TYO data center reaches 80% usage, we will stop the new order for TYO.Lite product until we have enough hardware.
Please be prepared if you have any needs in a short time.
We do not recommend you to hold many servers. For fair usage, DMIT keeps the right to stop services for you.
Our team decided to buy new servers after 2021.
As we estimate the price will be normal between Feb ~ Mar 2022
Before that, once our TYO data center reaches 80% usage, we will stop the new order for TYO.Lite product until we have enough hardware.
Please be prepared if you have any needs in a short time.
We do not recommend you to hold many servers. For fair usage, DMIT keeps the right to stop services for you.
NTT report an outage which might lead to high packet loss and/or latency increase.
Node HKG-0380L5W keeps getting high environment temperature warnings.
Our team has already asked IDC to check the situation.
Since the warning shows the temperature is higher than 100 degrees Celsius.
After figuring out where the problem is, we will decide to disable the temperature shut down temporarily, replace the modules later.
(or ask IDC to evacuate all people. xD)
Our team has already asked IDC to check the situation.
Since the warning shows the temperature is higher than 100 degrees Celsius.
After figuring out where the problem is, we will decide to disable the temperature shut down temporarily, replace the modules later.
(or ask IDC to evacuate all people. xD)
There is a major fiber cut between CoreSite LA1 and CoreSite LA2.
This led to DMIT CoreSite in-campus backbone failure.
By this, DMIT Los Angles network performance will be temporarily impacted.
Log from CoreSite NOC
Message Time Zone: EST.
- Feb 16 8:51PM
We are currently investigating a possible Fiber Cut between our LA1 and LA2 facility. Our Dark Fiber Vendor has been requested for a dispatch on priority. NO ETA as of yet
- Feb 17 12:21AM
A technician from our provider arrived onsite and confirmed the location of the fiber cut. This fiber is underground and will need a construction crew to come and dig up the street to confirm the extent of the damage. The construction/splicing team has been contacted according to our provider however, they have not provided a timeframe on when they will be onsite. We are escalating this issue and will provide updates as they become available.
- Feb 17 12:46AM
Construction and splice crews have been dispatched to the fiber cut site. ETA is 11:30 PM Pacific.
- Feb 17 2:05AM
Please be advised the construction and splicing crew has arrived onsite. They are currently breaking up the concrete and pulling the cable. Once that is complete they will start repairing/splicing the new cable.
Feb 17 5:35AM
Construction crew has completed pulling the new fiber and will start splicing now. Approximate restoration is 12:00pm Pacific. We will continue to provide updates as they become available.
Feb 17 10:05AM
Splice crews are now 70 percent complete with the splicing. Services my be seen restored while the splice crews work to complete the span. We will continue to provide updates as they become available or upon completion.
This led to DMIT CoreSite in-campus backbone failure.
By this, DMIT Los Angles network performance will be temporarily impacted.
Log from CoreSite NOC
Message Time Zone: EST.
- Feb 16 8:51PM
We are currently investigating a possible Fiber Cut between our LA1 and LA2 facility. Our Dark Fiber Vendor has been requested for a dispatch on priority. NO ETA as of yet
- Feb 17 12:21AM
A technician from our provider arrived onsite and confirmed the location of the fiber cut. This fiber is underground and will need a construction crew to come and dig up the street to confirm the extent of the damage. The construction/splicing team has been contacted according to our provider however, they have not provided a timeframe on when they will be onsite. We are escalating this issue and will provide updates as they become available.
- Feb 17 12:46AM
Construction and splice crews have been dispatched to the fiber cut site. ETA is 11:30 PM Pacific.
- Feb 17 2:05AM
Please be advised the construction and splicing crew has arrived onsite. They are currently breaking up the concrete and pulling the cable. Once that is complete they will start repairing/splicing the new cable.
Feb 17 5:35AM
Construction crew has completed pulling the new fiber and will start splicing now. Approximate restoration is 12:00pm Pacific. We will continue to provide updates as they become available.
Feb 17 10:05AM
Splice crews are now 70 percent complete with the splicing. Services my be seen restored while the splice crews work to complete the span. We will continue to provide updates as they become available or upon completion.
/var/log/DMIT-NOC.log
There is a major fiber cut between CoreSite LA1 and CoreSite LA2. This led to DMIT CoreSite in-campus backbone failure. By this, DMIT Los Angles network performance will be temporarily impacted. Log from CoreSite NOC Message Time Zone: EST. - Feb 16…
Dear Valued Customer,
Please see the below incident notification:
Update: Test is now completed and service should be restored. Please check your connectivity and confirm. If there are any issue you are experiencing please open a trouble ticket or if a ticket is already open please leave a comment through the ticket. Our local team members can look into your service further. This will now be closed.
Please see the below incident notification:
Update: Test is now completed and service should be restored. Please check your connectivity and confirm. If there are any issue you are experiencing please open a trouble ticket or if a ticket is already open please leave a comment through the ticket. Our local team members can look into your service further. This will now be closed.